# Multi-Modal Data-Efficient 3D Scene Understanding for Autonomous Driving

### v1

@@ 1 Introduction
Building on this premise, this work introduces LaserMix++, an advanced data-efficient 3D scene understanding framework that expands the semi-supervised learning paradigm to incorporate multi-modal data integration. The baseline LaserMix framework [10] leveraged spatial priors inherent in LiDAR data (as shown in Figure 1(a)) by mixing laser beams from different scans to enhance the consistency and confidence of predictions across unlabeled datasets. This approach utilized the geometric distribution of LiDAR-acquired driving scenes to infer the scene semantics with minimal supervision, setting a seminar yet strong benchmark in semi-supervised 3D scene understanding [10].
Expanding upon this, LaserMix++ integrates additional sensor modalities, specifically camera data, to address the complexities of autonomous driving environments more comprehensively. By harnessing both LiDAR and camera inputs, LaserMix++ aims to exploit the complementary nature of spatial priors from LiDAR and textural details from camera images. This multi-modal approach enhances the model’s ability to generalize across different scenarios, particularly in conditions where supervision signals are not sufficiently available [10, 11, 7]. LaserMix++ introduces three novel components to achieve better multi-modal integration:
@@
  • Camera-to-LiDAR Feature Distillation: Leveraging recent endeavors in image segmentation, we propose to extract semantically rich features from camera images and integrate them into the LiDAR data processing stream. This process helps in enhancing the data-efficient learning of the LiDAR point cloud data, particularly by improving the feature representation in environments where LiDAR data annotation alone is insufficient.
@@
The combination of the above components not only addresses the limitations of using single-modality data but also boosts the robustness and accuracy of the semi-supervised LiDAR segmentation model. Each step in the expansion of our LaserMix++ framework builds logically on the last, ensuring that the enhancements contribute meaningfully to the overall effectiveness of the multi-modal data-efficient 3D scene understanding system.
@@ 3.2 LaserMix
where $x_{i}^{a_{j}}$ is the data crop of $x_{i}$ confined within area area $a_{j}$. Correspondingly, the semantic labels are mixed in the same way as Equation 7. It is worth highlighting that LaserMix is directly applied to the LiDAR point clouds and is thus agnostic to the various LiDAR representations, such as range view [40], bird’s eye view [37], raw points [89], sparse voxel [33], and multi-view fusion [41] representations. This versatility allows LaserMix to be effectively integrated across a broad spectrum of existing LiDAR segmentation frameworks, without necessitating any modifications to the underlying data or network structures.
@@
$\mathcal{L}_{\text{mt}}=||\mathcal{G}_{\theta}^{s}(x)-\mathcal{G}_{\theta}^{t}(x)||^{2}_{2}\,,$ (9)
@@ 4 LaserMix++
Assuming that the driving perception system consists of one LiDAR and one camera sensor [^1] This simple configuration can be easily extended to single-LiDAR multi-camera and multi-LiDAR multi-camera scenarios. that have been well calibrated and synchronized, we can obtain, at a certain data acquisition frequency, a pair of LiDAR point cloud $x_{\text{p}}=(p^{x},p^{y},p^{z})$ and camera image $x_{\text{img}}$. To establish 2D-3D correspondences for a given pair $\{x_{\text{p}},x_{\text{img}}\}$, we first project point cloud $x_{\text{lidar}}$ onto the image plane $(p^{u},p^{v})$ based on sensor calibration parameters, i.e.,
@@
where $\Gamma_{K}$ denotes the camera intrinsic matrix and $\Gamma_{c\leftarrow l}$ is the transformation matrix from the LiDAR to the camera. It is worth mentioning that the correspondences between points and pixels do not always hold due to the possible miss-overlap between the field-of-views of the LiDAR and the camera. We create a correspondence mask $\mathcal{M}$ of the same length with $x_{\text{p}}$ to query valid point-pixel relationships. For LiDAR points that do not have a corresponding pixel, we pad their pixel correspondences with zero.
@@ 4.2 Multi-Modal LaserMix
The area set $\{a_{1},a_{2},...,a_{m}\}$ can be obtained in the same way as Sec. 3.2. The multi-modal variant of LaserMix leverages off-the-shelf visual information from synchronized camera images to assist the consistency regularization in Sec. 3.3. Such an approach further enhances the effect of spatial priors in driving scenes, since the LiDAR points have been “painted” with extra texture information from the images. After obtaining mixed LiDAR-image pairs, we propose the following two operations in the LaserMix++ framework to assist the data-efficient 3D scene understanding task with such multi-modal inputs.
@@ 4.3 Camera-to-LiDAR Feature Distillation
Over the past few years, image segmentation models have become both more efficient and highly accurate on large-scale benchmarks [91, 92, 93, 94, 2]. Most recent models are trained across a wide spectrum of image collections and, as a return, show promising zero-shot generalizability to unseen domains [95, 96, 97]. Such a versatile capability opens up new possibilities for driving perceptions from RGB cameras [4]. Observing the promising image segmentation capability from these foundation models, we propose to leverage LiDAR-camera correspondences for data-efficient 3D scene understanding, without using ground truth image labels. Specifically, given $\{x_{\text{p}},x_{\text{img}}\}$ pairs, we extract the point cloud and image features from the backbone of the Student network and an image segmentation backbone $\hat{\mathcal{G}}_{\xi}^{\text{img}}$ with a pretrained parameter set ξ as follows:
@@
where $\hat{\mathcal{G}}_{\hat{\theta}}$ with parameters $\hat{\theta}$ is the backbone of the Student network $\mathcal{G}_{\theta}^{t}$, and $\mathcal{H}_{\zeta}^{s}$ with parameters $\zeta$ is a projection layer used to map the dimension of the point cloud feature to that of the image feature. According to the known sensor calibration parameters in Equation 10, we obtain the paired point-image features $\{F_{\text{p}},\tilde{F}_{\text{img}}\}$, where $\tilde{F}_{\text{img}}$ has been aligned to match the shape with $F_{\text{p}}$. To distill the knowledge from $\hat{\mathcal{G}}_{\xi}^{\text{img}}$ to $\hat{\mathcal{G}}_{\hat{\theta}}$, we design the camera-to-LiDAR feature distillation loss $\mathcal{L}_{\text{c2l}}$, which aims to minimize the cosine distance between two features as follows:
$\mathcal{L}_{\text{c2l}}=\mathcal{M}\odot(1-\langle F_{\text{p}}\;,\;\tilde{F}_{\text{img}}\rangle)\,,$ (13)
where $\odot$ denotes the element-wise multiplication; $\langle$,$\rangle$ , ⟩ calculates the inner product of two features. This cross-sensor optimization objective transfers the semantically more coherent features from RGB images to LiDAR point clouds, which assists the LiDAR segmentation model in learning better representation under data-efficient scenarios.
@@ 4.4 Language-Driven Knowledge Guidance
The most prevailing vision-language models like CLIP [39] have been widely exploited in the image domain to enable open-vocabulary predictions. Such a paradigm enables a more holistic perception beyond the fixed label set, which is especially suitable for driving scenes. Recent endeavors on open-vocabulary image segmentation attempt to train large-scale multi-task models that align well with text inputs [98, 96, 97]. To enrich the supervision signals for data-efficient 3D scene understanding, we propose to use these pretrained models along with open vocabularies to generate auxiliary labels on unlabeled LiDAR point clouds. Specifically, given the class names as text prompts, we first use the CLIP’s text encoder $\mathcal{G}^{\text{txt}}_{\varrho}$ to extract text embedding $F^{{}^{\prime}}_{\text{txt}}$, with $\varrho$ being the pretrained parameters. Next, we feed the image features $F_{\text{img}}$ extracted from Equation 12 to the pretrained open-vocabulary image segmentation head $\mathcal{H}_{\varsigma}^{\text{img}}$ and also the point cloud to the LiDAR segmentation model $\mathcal{G}_{\theta}^{s}$. The whole process can be described as follows:
$\begin{split}F^{{}^{\prime}}_{\text{p}}=\mathcal{G}_{\theta}^{s}(x_{\text{p}})\,,~{}~{}~{}F^{{}^{\prime}}_{\text{img}}=\langle\mathcal{H}_{\varsigma}^{\text{img}}(F_{\text{img}})\;,\;F^{{}^{\prime}}_{\text{txt}}\rangle\,,\end{split}$ (14)
where $\varsigma$ denotes the parameters of $\mathcal{H}_{\varsigma}^{\text{img}}$. $F^{{}^{\prime}}_{\text{p}}$ and $F^{{}^{\prime}}_{\text{img}}$ are the non-probabilistic outputs from the models. Similar to Equation 13, we first find the paired point-image outputs $\{F^{{}^{\prime}}_{\text{p}},\tilde{F}^{{}^{\prime}}_{\text{img}}\}$. To achieve language-driven knowledge guidance from images to LiDAR point clouds, we minimize the cosine distance between two outputs as follows:
$\mathcal{L}_{\text{lkg}}=\mathcal{M}\odot(1-\langle F^{{}^{\prime}}_{\text{p}}\;,\;\tilde{F}^{{}^{\prime}}_{\text{img}}\rangle)\,.$ (15)
@@ 5 Experiments
5.1 Datasets
5.2 Experimental Setups
Implementation Details. Our LaserMix++ framework is established based on the MMDetection3D codebase [106]. We denote the supervised-only baseline as sup.-only in our experiments. For semi-supervised learning, the labeled/unlabeled data are selected in four ways as shown in Figure 4: 1) random sampling, 2) uniform sampling, 3) sequential sampling, and 4) ST-RFD [11] sampling strategies. Due to the lack of previous works, we also compare consistency regularization [18, 21, 61] and entropy minimization [67] methods from semi-supervised image segmentation. The number of spatial area sets $m$ in multi-modal LaserMix is uniformly sampled from 2 to 6. The loss weights $\lambda_{\text{mix}}$, $\lambda_{\text{mt}}$, $\lambda_{\text{c2l}}$, and $\lambda_{\text{lkg}}$ in Equation 16 are set to 2.0, 250, 1.5, and 1.0, respectively. All experiments are implemented using PyTorch on eight NVIDIA A100 GPUs. All baseline models are trained with a batch size of 2 on each GPU, along with the AdamW optimizer [107], OneCycle learning rate schedule [108], and a learning rate of 0.008. We do not include any type of test-time augmentation or model ensemble during the evaluation.
@@ 5.2 Experimental Setups
c $43.2$ $\mathbf{+5.7}$ $57.1$ $\mathbf{+4.0}$ $58.3$ $\mathbf{+2.2}$ $59.8$ $\mathbf{+2.4}$ $45.6$ $\mathbf{+3.5}$ $64.3$ $\mathbf{+3.9}$ $67.8$ $\mathbf{+2.4}$ $71.6$ $\mathbf{+2.2}$
@@
d $46.5$ $\mathbf{+9.0}$ $59.3$ $\mathbf{+6.2}$ $60.4$ $\mathbf{+4.3}$ $61.9$ $\mathbf{+4.5}$ $46.0$ $\mathbf{+3.9}$ $64.1$ $\mathbf{+3.7}$ $69.5$ $\mathbf{+4.1}$ $72.3$ $\mathbf{+2.9}$
@@
e $48.3$ $\mathbf{+0.9}$ $60.7$ $\mathbf{+0.6}$ $61.4$ $\mathbf{+0.4}$ $63.1$ $\mathbf{+0.5}$ $50.1$ $\mathbf{+0.6}$ $68.8$ $\mathbf{+0.6}$ $71.0$ $\mathbf{+0.4}$ $73.1$ $\mathbf{+0.1}$
@@
f $49.6$ $\mathbf{+2.2}$ $61.2$ $\mathbf{+1.1}$ $61.9$ $\mathbf{+0.9}$ $63.4$ $\mathbf{+0.8}$ $50.8$ $\mathbf{+1.3}$ $69.4$ $\mathbf{+1.2}$ $71.3$ $\mathbf{+0.7}$ $73.3$ $\mathbf{+0.3}$
@@
g $\mathbf{50.1}$ $\mathbf{+2.7}$ $\mathbf{61.9}$ $\mathbf{+1.8}$ $\mathbf{62.4}$ $\mathbf{+1.4}$ $\mathbf{63.7}$ $\mathbf{+1.1}$ $\mathbf{51.6}$ $\mathbf{+2.1}$ $\mathbf{69.8}$ $\mathbf{+1.6}$ $\mathbf{71.7}$ $\mathbf{+1.1}$ $\mathbf{73.7}$ $\mathbf{+0.7}$
5.3 Comparative Study
Generalize to Image Segmentation. The proposed scene prior-based consistency regularization can be extended to image domains since they also reflect real-world driving scenarios. To validate this, we conduct experiments on Cityscapes [94] and show the results in Tab. V. As can be seen, our approaches achieved improved performance over the strong image segmentation baselines [18, 21, 19, 20]. Such a generalizability ensures our approaches to be universally applicable to different sensor modalities.

### v2

@@ 1 Introduction
Building on this premise, this work introduces LaserMix++, an advanced data-efficient 3D scene understanding framework that expands the semi-supervised learning paradigm to incorporate multi-modal data integration. The baseline LaserMix [10] is a single-modal framework that leverages spatial priors inherent in LiDAR data (see Figure 1(a)) by mixing laser beams from LiDAR scans to enhance the consistency and confidence of predictions across unlabeled datasets. This approach utilized the geometric distribution of LiDAR-acquired driving scenes to infer the scene semantics with minimal supervision, setting a seminar yet strong benchmark in 3D scene understanding [10].
Expanding upon this foundation, LaserMix++ integrates additional sensor modalities, specifically camera data, to address the complexities of autonomous driving environments more comprehensively. While the original LaserMix [10] focused on utilizing spatial continuity within LiDAR data to improve semantic understanding, LaserMix++ builds on this by incorporating textural and contextual information from camera images. By harnessing both LiDAR and camera inputs, LaserMix++ aims to exploit the complementary nature of spatial priors from LiDAR and textural details from camera images. This multi-modal approach enhances the model’s ability to generalize across different scenarios, particularly in conditions where supervision signals are not sufficiently available [10, 11, 7]. LaserMix++ introduces three novel components to achieve better multi-modal integration:
@@
  • Camera-to-LiDAR Feature Distillation: Leveraging recent endeavors in image segmentation, we propose to extract semantically rich features from camera images and integrate them into the LiDAR data processing stream. This method aims to effectively bridge the gap between 2D image data and 3D point clouds, aligning with the core principle of LaserMix [10] by enhancing the scene consistency across different modalities. This process helps in enhancing the data-efficient learning of the LiDAR point cloud data, particularly by improving the feature representation in environments where LiDAR data annotation alone is insufficient.
@@
By integrating these multi-modal components, LaserMix++ retains the strengths of the original LaserMix [10] – such as leveraging spatial priors from LiDAR – while adding the ability to incorporate and align additional sensor data. This combination not only addresses the limitations of using single-modality data but also boosts the robustness and accuracy of the semi-supervised LiDAR segmentation model. Each step in the expansion of our LaserMix++ framework builds logically on the last, ensuring that the enhancements contribute meaningfully to the overall effectiveness of the multi-modal data-efficient 3D scene understanding system.
@@ 3.2 LaserMix
where $x_{i}^{a_{j}}$ is the data crop of $x_{i}$ confined within area $a_{j}$. Correspondingly, the semantic labels are mixed in the same way as Equation 7. It is worth highlighting that LaserMix is directly applied to the LiDAR point clouds and is thus agnostic to the various LiDAR representations, such as range view [40], bird’s eye view [37], raw points [89], sparse voxel [33], and multi-view fusion [41] representations. This versatility allows LaserMix to be effectively integrated across a broad spectrum of existing LiDAR segmentation frameworks, without necessitating any modifications to the underlying data or network structures. While LaserMix partitions the LiDAR point cloud for enhancing feature learning, concerns may arise regarding the integrity of object instances, such as cars, when these are split across different segments. However, given the large spatial area typically involved, the likelihood of significant object partitioning is low. Moreover, LaserMix’s use of both ground truth and high-confidence pseudo labels provides robust supervision, ensuring accurate object interpretation even when partitioning occurs. This design allows LaserMix to leverage the benefits of data augmentation without compromising object integrity, thereby enhancing the robustness and generalization of the model.
@@
$\mathcal{L}_{\text{mt}}=\frac{1}{|x|}\sum_{i=1}^{|x|}||\mathcal{G}_{\theta}^{s,(i)}(x)-\mathcal{G}_{\theta}^{t,(i)}(x)||^{2}_{2}\,,$ (9)
@@ 4 LaserMix++
Assuming that the driving perception system consists of one LiDAR and one camera sensor [^1] This simple configuration can be easily extended to both single-LiDAR multi-camera and multi-LiDAR multi-camera scenarios. that have been well calibrated and synchronized, we can obtain, at a certain data acquisition frequency, a pair of LiDAR point cloud $x_{\text{p}}=(p^{x},p^{y},p^{z})$ and camera image $x_{\text{img}}$. To establish 2D-3D correspondences for a given pair $\{x_{\text{p}},x_{\text{img}}\}$, we first project point cloud $x_{\text{p}}$ onto the image plane $(p^{u},p^{v})$ based on sensor calibration parameters:
@@
where $\Gamma_{K}$ denotes the camera intrinsic matrix and $\Gamma_{c\leftarrow l}$ is the transformation matrix from the LiDAR to the camera.
It is worth mentioning that the correspondences between points and pixels do not always hold due to the possible mis-overlap between the field-of-views of the LiDAR and the camera. To handle this, we generate a correspondence mask $\mathcal{M}$ to establish valid point-pixel relationships. The process begins by projecting LiDAR points onto the image plane using the known extrinsic and intrinsic camera parameters, allowing us to determine which LiDAR points have corresponding pixels in the image. If a LiDAR point $x_{\text{p}}$ projects onto a valid pixel location, the corresponding entry in the mask $\mathcal{M}$ is set to $1$. For LiDAR points that project outside the image boundaries or onto invalid regions, the corresponding entry in $\mathcal{M}$ is set to 00. These entries are then padded with zeros in the pixel correspondence data. This automated process leverages existing geometric transformations and does not require additional manual data annotation, ensuring scalability and efficiency.
@@ 4.2 Multi-Modal LaserMix
The area set $\{a_{1},a_{2},...,a_{m}\}$ can be obtained in the same way as Sec. 3.2. The multi-modal variant of LaserMix leverages off-the-shelf visual information from synchronized camera images to assist the consistency regularization in Sec. 3.3. Such an approach further enhances the effect of spatial priors in driving scenes, since the LiDAR points have been “painted” with extra texture information from the images. The use of the correspondence mask $\mathcal{M}$ ensures that only valid LiDAR point-pixel pairs are utilized, maintaining the consistency of the multi-modal information. After obtaining mixed LiDAR-image pairs, we propose the following two operations in the LaserMix++ framework to assist the data-efficient 3D scene understanding task with such multi-modal inputs.
@@ 4.3 Camera-to-LiDAR Feature Distillation
Over the past few years, image segmentation models have become both more efficient and highly accurate on large-scale benchmarks [91, 92, 93, 94, 2]. Most recent models are trained across a wide spectrum of image collections and, as a result, demonstrate promising zero-shot generalizability to unseen domains [95, 96, 97]. This versatile capability opens up new possibilities for driving perceptions from RGB cameras [4]. Leveraging these advances, we aim to transfer the semantic richness of pre-trained image models to LiDAR-based 3D scene understanding in a data-efficient manner, avoiding the need for ground truth image labels.
Specifically, given $\{x_{\text{p}},x_{\text{img}}\}$ pairs, we extract the point cloud and image features using the Student network’s backbone and a pre-trained image segmentation backbone $\hat{\mathcal{G}}_{\xi}^{\text{img}}$, parameterized by $\xi$:
@@
where $\hat{\mathcal{G}}_{\hat{\theta}}$ with parameters $\hat{\theta}$ is the backbone of the Student network $\mathcal{G}_{\theta}^{t}$, and $\mathcal{H}_{\zeta}^{s}$ is a projection layer mapping the dimension of point cloud features to match that of image features. Based on the sensor calibration parameters from Equation 10, we align the features to create paired point-image features $\{F_{\text{p}},\tilde{F}_{\text{img}}\}$. A key aspect of our approach is ensuring that the integration of image-based semantics does not diminish the unique spatial features of the point cloud.
To achieve this goal, we introduce a feature alignment strategy augmented with residual connections. This design allows the network to enhance point cloud features with semantic information from the image while maintaining the point cloud’s inherent geometrical properties. We then define the camera-to-LiDAR feature distillation loss $\mathcal{L}_{\text{c2l}}$ to minimize the cosine distance between the paired features:
$\mathcal{L}_{\text{c2l}}=\frac{1}{\sum_{i=1}^{|x_{\text{p}}|}\mathcal{M}^{(i)}}\sum_{i=1}^{|x_{\text{p}}|}\left(\mathcal{M}^{(i)}\cdot\left(1-\langle F_{\text{p}}^{(i)},\tilde{F}_{\text{img}}^{(i)}\rangle\right)\right),$ (13)
where $\langle$,$\rangle$ , ⟩ calculates the inner product. The use of cosine distance ensures that the semantic alignment process enriches the point cloud features without overwhelming them, thereby achieving a balanced integration of both semantic and geometric information. This approach facilitates the transfer of semantically coherent features from images to LiDAR, enhancing the LiDAR segmentation model’s performance in a data-efficient manner.
@@ 4.4 Language-Driven Knowledge Guidance
Vision-language models, such as CLIP [39], have demonstrated significant success in enabling open-vocabulary predictions by aligning visual and textual information. This capability is highly beneficial for driving scenarios, where a broader understanding beyond fixed label sets is required. Recent work in open-vocabulary image segmentation has further leveraged these models to train large-scale multi-task networks that can effectively align with textual inputs [98, 96, 97]. In our framework, we harness these capabilities to provide enriched supervision signals for 3D scene understanding using unlabeled LiDAR point clouds.
Given class names as text prompts, we use the CLIP text encoder $\mathcal{G}^{\text{txt}}_{\varrho}$ to extract text embeddings $F^{{}^{\prime}}_{\text{txt}}$, parameterized by $\varrho$. These embeddings, combined with the image features $F_{\text{img}}$ from Equation 12, are processed through the pre-trained open-vocabulary image segmentation head $\mathcal{H}_{\varsigma}^{\text{img}}$ and the LiDAR segmentation model $\mathcal{G}_{\theta}^{s}$:
$\begin{split}F^{{}^{\prime}}_{\text{p}}=\mathcal{G}_{\theta}^{s}(x_{\text{p}})\,,~{}~{}~{}F^{{\prime}}_{\text{img}}=\mathcal{H}{\varsigma}^{\text{img}}(F_{\text{img}})\circ F^{{\prime}}_{\text{txt}}\,,\end{split}$ (14)
where $\varsigma$ represents the parameters of $\mathcal{H}_{\varsigma}^{\text{img}}$. $F^{{}^{\prime}}_{\text{p}}$ and $F^{{}^{\prime}}_{\text{img}}$ denote non-probabilistic outputs from the models, which are then paired as $\{F^{{}^{\prime}}_{\text{p}},\tilde{F}^{{}^{\prime}}_{\text{img}}\}$. Symbol $\circ$ represents the alignment operation that combines the image feature vector $\mathcal{H}{\varsigma}^{\text{img}}(F_{\text{img}})$ with the text feature $F^{\prime}_{\text{txt}}$. In our actual implementation, we use the text and image encoders from CLIP [39] to extract the text embedding and image embedding, respectively. We then use a cosine similarity loss to align the text and image features, resulting in text-aligned $F^{{}^{\prime}}_{\text{img}}$.
To integrate semantic cues derived from text into the point cloud domain effectively, we employ a weighted feature fusion approach that respects the spatial integrity of the LiDAR data. This strategy, along with residual connections, ensures that the semantic information enhances the understanding of the scene without compromising the distinct structural characteristics of the point cloud. By maintaining this balance, our approach provides a richer and more robust understanding of the 3D environment.
To realize the language-driven knowledge guidance objective, we minimize the cosine distance between the point features and the text-aligned image features, that is:
$\mathcal{L}_{\text{lkg}}=\frac{1}{\sum_{i=1}^{|x_{\text{p}}|}\mathcal{M}^{(i)}}\sum_{i=1}^{|x_{\text{p}}|}\mathcal{M}^{(i)}\cdot\left(1-\langle F_{\text{p}}^{\prime,(i)},\tilde{F}^{\prime,(i)}_{\text{img}}\rangle\right)\,.$ (15)
This loss function not only aligns the features but also ensures that the point cloud’s unique spatial features are preserved during the distillation process, promoting a harmonious integration of multi-modal information.
@@ 5 Experiments
c $40.0$ $\mathbf{+2.5}$ $54.7$ $\mathbf{+1.6}$ $56.6$ $\mathbf{+0.5}$ $57.9$ $\mathbf{+0.5}$ $43.9$ $\mathbf{+1.8}$ $61.6$ $\mathbf{+1.2}$ $66.2$ $\mathbf{+0.8}$ $69.9$ $\mathbf{+0.5}$
$41.2$ $\mathbf{+3.7}$ $55.3$ $\mathbf{+2.2}$ $56.9$ $\mathbf{+0.8}$ $58.2$ $\mathbf{+0.8}$ $44.5$ $\mathbf{+2.4}$ $61.8$ $\mathbf{+1.4}$ $66.2$ $\mathbf{+0.8}$ $69.6$ $\mathbf{+0.2}$
d $43.2$ $\mathbf{+5.7}$ $57.1$ $\mathbf{+4.0}$ $58.3$ $\mathbf{+2.2}$ $59.8$ $\mathbf{+2.4}$ $45.6$ $\mathbf{+3.5}$ $64.3$ $\mathbf{+3.9}$ $67.8$ $\mathbf{+2.4}$ $71.6$ $\mathbf{+2.2}$
@@
e $46.5$ $\mathbf{+9.0}$ $59.3$ $\mathbf{+6.2}$ $60.4$ $\mathbf{+4.3}$ $61.9$ $\mathbf{+4.5}$ $46.0$ $\mathbf{+3.9}$ $64.1$ $\mathbf{+3.7}$ $69.5$ $\mathbf{+4.1}$ $72.3$ $\mathbf{+2.9}$
@@
f $48.3$ $\mathbf{+0.9}$ $60.7$ $\mathbf{+0.6}$ $61.4$ $\mathbf{+0.4}$ $63.1$ $\mathbf{+0.5}$ $50.1$ $\mathbf{+0.6}$ $68.8$ $\mathbf{+0.6}$ $71.0$ $\mathbf{+0.4}$ $73.1$ $\mathbf{+0.1}$
@@
g $49.6$ $\mathbf{+2.2}$ $61.2$ $\mathbf{+1.1}$ $61.9$ $\mathbf{+0.9}$ $63.4$ $\mathbf{+0.8}$ $50.8$ $\mathbf{+1.3}$ $69.4$ $\mathbf{+1.2}$ $71.3$ $\mathbf{+0.7}$ $73.3$ $\mathbf{+0.3}$
@@
h $\mathbf{50.1}$ $\mathbf{+2.7}$ $\mathbf{61.9}$ $\mathbf{+1.8}$ $\mathbf{62.4}$ $\mathbf{+1.4}$ $\mathbf{63.7}$ $\mathbf{+1.1}$ $\mathbf{51.6}$ $\mathbf{+2.1}$ $\mathbf{69.8}$ $\mathbf{+1.6}$ $\mathbf{71.7}$ $\mathbf{+1.1}$ $\mathbf{73.7}$ $\mathbf{+0.7}$
@@
[…]
Method Mix KITTI nuScenes Scribble
mIoU mAcc mIoU mAcc mIoU mAcc
MinkUNet [52] $66.9$ $92.4$ $76.4$ $94.1$ $61.2$ $88.5$
$\mathbf{68.1}$ $\mathbf{92.6}$ $\mathbf{76.9}$ $\mathbf{94.3}$ $\mathbf{62.0}$ $\mathbf{89.0}$
Cylinder3D [33] $63.7$ $91.0$ $75.8$ $93.7$ $58.8$ $87.4$
$\mathbf{64.7}$ $\mathbf{91.3}$ $\mathbf{78.1}$ $\mathbf{94.0}$ $\mathbf{59.5}$ $\mathbf{88.6}$
PolarNet [37] $57.2$ $91.0$ $71.7$ $93.1$ $55.7$ $87.6$
$\mathbf{59.6}$ $\mathbf{91.2}$ $\mathbf{72.1}$ $\mathbf{93.2}$ $\mathbf{56.5}$ $\mathbf{88.1}$
FRNet [50] $64.1$ $92.2$ $76.8$ $93.4$ $57.6$ $88.3$
$\mathbf{66.4}$ $\mathbf{92.4}$ $\mathbf{77.6}$ $\mathbf{93.9}$ $\mathbf{59.6}$ $\mathbf{88.8}$
5.1 Datasets
5.2 Experimental Setups
Implementation Details. Our LaserMix++ framework is established based on the MMDetection3D codebase [107]. We denote the supervised-only baseline as sup.-only in our experiments. For semi-supervised learning, the labeled/unlabeled data are selected in four ways as shown in Figure 4: 1) random sampling, 2) uniform sampling, 3) sequential sampling, and 4) ST-RFD [11] sampling strategies. Due to the lack of previous works, we also compare consistency regularization [18, 21, 61] and entropy minimization [67] methods from semi-supervised image segmentation. The number of spatial area sets $m$ in multi-modal LaserMix is uniformly sampled from 2 to 6. The loss weights $\lambda_{\text{mix}}$, $\lambda_{\text{mt}}$, $\lambda_{\text{c2l}}$, and $\lambda_{\text{lkg}}$ in Equation 16 are set to 2.0, 250, 1.5, and 1.0, respectively. All experiments are implemented using PyTorch on eight NVIDIA A100 GPUs. All baseline models are trained with a batch size of 2 on each GPU, along with the AdamW optimizer [108], OneCycle learning rate schedule [109], and a learning rate of 0.008. We do not include any type of test-time augmentation or model ensemble during the evaluation. For additional details, kindly refer to the Appendix.
5.3 Comparative Study
Generalize to Image Segmentation. The proposed scene prior-based consistency regularization can be extended to image domains since they also reflect real-world driving scenarios. To validate this, we conduct experiments on Cityscapes [94] and show the results in Tab. V. As can be seen, our approaches achieved improved performance over the strong image segmentation baselines [18, 21, 19, 20]. Such a generalizability ensures our approaches are universally applicable to different sensor modalities.